# Position: Membership Inference Attack Should Move On to Distributional Statistics for Distilled Generative Models

### v3

@@ 1 Introduction
The success of I-MIAs mainly relies on a phenomenon in training such large-scale generative models: well-trained models tend to overfit to their training set111 Note that, for I-MIAs used in classifiers, overfitting is not the only way to help distinguish between members and non-members. For example, Carlini et al. (2022) successfully identified member data by analyzing model behavior signals, such as the output confidence distribution and gradient norms, even when a classifier does not overfit the data. Moreover, it does not mean that the overfitting is a golden rule that always makes I-MIAs successful. When the large-scale models are trained with extremely large datasets (e.g., ChatGPT), I-MIAs could be meaningless (Zhang et al., 2024). We will discuss this in detail in Section 3.1., exhibiting different behaviors between training instances (i.e., called members) and test instances (i.e., called non-members) (Yeom et al., 2018), which is also known as the memorization effect (Yeom et al., 2018). Based on this phenomenon, many I-MIAs consider using training objectives as criteria to distinguish between members and non-members. For example, for diffusion models, Duan et al. (2023) use estimation errors during the denoising process as the member-detection criterion; for large language models, Ye et al. (2024) and Shejwalkar and Houmansadr (2021) use log-likelihood for the same purpose. Recent studies (Carlini et al., 2023; Hu and Pang, 2023; Duan et al., 2023) have investigated these criteria extensively, hypothesizing that generative models can memorize their training instances.
@@
The failure of I-MIA raises a question: Should we discard the distilled generative models when we care about privacy? The answer is negative, as we discover a distribution-based memory chain between a student-teacher pair: student-generated data exhibit a significantly stronger distributional alignment with teacher’s members than non-members, making it possible to determine if a student model has knowledge from unauthorized data. Specifically, in Sec. 4.1, we use maximum mean discrepancy (MMD) to compare student-generated samples with both teacher’s member and non-member data distributions (Fig. 3(c)). Through repeated random subset samplings, we observe a consistent statistical pattern–distances to non-member data concentrate at higher values than to member data, suggesting that the student preserves statistical signals exhibiting stronger alignment with the teacher’s member distribution than non-member distributions, despite the failure of I-MIAs. Thus, to reliably audit the privacy violation of distilled generative models, we posit the following statement in the field of generative models:
@@ 2 Preliminary: MIAs for Generative Models
We first introduce some basic concepts and relevant literature in this section. For readers who are familiar with MIAs for generative models, it is safe to skip this section. MIAs evaluate whether a specific instance is used during model training. Let ${\mathcal{X}}$ be a data space and ${\mathcal{D}}^{\rm mem}\subset{\mathcal{X}}$ be a member set used to train a generative model $G:{\mathcal{Z}}\to{\mathcal{X}}$ that transforms noises sampled from a latent distribution ${\mathbf{z}}\sim p({\mathbf{z}})$ into synthetic data ${\mathbf{x}}=f_{g}({\mathbf{z}})\in{\mathcal{X}}$. Given a query sample ${\bm{x}}_{q}\in{\mathcal{X}}$, an MIA constructs a binary classifier ${\mathcal{A}}:{\mathcal{X}}\times G\to\{0,1\}$ that predicts the membership attribution of ${\mathbf{x}}_{q}$ as ${\mathcal{A}}({\bm{x}}_{q},G)=1$ if ${\bm{x}}_{q}\in{\mathcal{D}}^{\rm mem}$, and 00 otherwise. Existing MIAs use instance-level scores to distinguish members from non-members, a strategy we term instance-level MIA (I-MIA), including reference-based and intrinsic-based methods that are introduced below.
@@ 2.1 Brief introduction for instance-level MIAs
where ${\mathcal{D}}^{\rm mem}_{i}$ denotes the training dataset of the $i$-th reference model. The membership inference decision is then based on the difference between the target model $\phi(G,{\bm{x}})$ and these groups. For example, ${\mathcal{A}}({\bm{x}},G)=1$ if a difference metric $\Delta({\bm{x}},G)>\tau$ and 00 otherwise, where $\Delta({\bm{x}},G)\triangleq\mathrm{sim}(\phi(G,{\bm{x}}),{\mathcal{M}}_{1})-\mathrm{sim}(\phi(G,{\bm{x}}),{\mathcal{M}}_{0})$, $\mathrm{sim}(\cdot,\cdot)$ is a function to measure the similarity between two models, and $\tau$ is the decision threshold. However, this approach faces practical infeasibility for computationally intensive generative models like diffusion models, as training $n\geq 1$ reference models would increase the attack cost.
@@ 2.2 How I-MIAs are applied to large-scale generative models?
Large-scale generative model in text domain-LLMs (Guo et al., 2025; Touvron et al., 2023) are auto-regressive transformers (Tang et al., 2024) trained on massive text corpora to model the log-likelihood of each next token given its preceding context. Several MIA strategies have been developed for LLMs. PCA (Ye et al., 2024) detects membership by comparing the target input with a synthetic version generated by word-swapping. A large log-likelihood gap indicates membership. The Min-K(Shokri et al., 2017) computes the average log-likelihood of the $k$ tokens with the least confidence in a given input. Typically, member data exhibit higher log-likelihoods for these tokens compared to non-members. Recall (Xie et al., 2024) introduces a non-member prefix to condition the model and quantifies the resulting change in log-likelihood. If the log-likelihood changes significantly, the data is inferred to be a member sample.
Large-scale generative model in vision domain-Diffusion models (Ho et al., 2020; Song et al., 2020), which operate through a forward process $q({\bm{x}}_{t}|{\bm{x}}_{0})$ that progressively adds Gaussian noise $\epsilon\sim{\mathcal{N}}(0,{\mathbf{I}})$ to the input ${\bm{x}}_{0}$ and a reverse noise removal process $p_{G}({\bm{x}}_{0}|{\bm{x}}_{t})$ that reconstructs the original ${\bm{x}}_{0}$ over time steps $t\in[0,T]$. Several attacks have been proposed in this context and have been evaluated on a class of diffusion models, the denoising diffusion probabilistic model (DDPM) (Ho et al., 2020). According to DDPM objective loss $\mathcal{L}_{\text{MSE}}=\left\|\epsilon_{G}(x_{t},t)-\epsilon\right\|^{2}$, SecMI (Duan et al., 2023) performs an attack by monitoring this loss across different time-steps $t$. ReDiffuse (Li et al., 2024) investigates reconstruction stability under noise perturbations, based on the observation that member data tend to yield more consistent reconstructions. GSA (Pang et al., 2023) examines gradient dynamics during model retraining and finds that member data typically induces smaller parameter updates.
@@ 3.1 Large-scale generative models challenge the effectiveness of I-MIAs
I-MIAs face challenges when attacking LLMs. Although I-MIA remains widely used for large-scale generative models, it is unreliable, specifically for LLMs (Dong et al., 2024; Ye et al., 2024). This is because extensive training on massive corpora and substantial model capacity of LLMs would lead to similar behaviors between individual members and non-members exploited by I-MIAs (Ye et al., 2024). When processing an input, LLMs consistently produce high-confidence outputs regardless of whether it is part of training data, reducing the discriminative power of instance-level metrics (Dong et al., 2024). Moreover, Zhang et al. (2024) argues that I-MIAs on LLMs suffer from unboundable false positive rates (FPR)—a critical metric for evaluating the validity of MIAs when used as evidence to allege the use of unauthorized data (Carlini et al., 2022; Zhang et al., 2024). The reason is that LLMs are trained with massive web-scale corpora, such that there are no ‘true’ non-members (as LLMs indirectly learn too much data (Zhang et al., 2024)).
I-MIAs face challenges when attacking diffusion models. In the case of generative models in the visual domain—diffusion models, existing MIAs have successfully identified members in non-distilled settings (Duan et al., 2023; Li et al., 2024). However, there are still issues for both reference-based MIAs and intrinsic-based MIAs. First, reference-based MIAs are impractical for diffusion models due to (1) the high cost of training a reference model (Ho et al., 2020), and (2) the impossibility of recovering the original input noise, which is sampled randomly during training (Ho et al., 2020; Kim et al., 2023).
@@
Second, several intrinsic-based MIAs have shown promising results on DDPMs (Ho et al., 2020). Since the DDPM objective (described in Sec. 2.2) encourages accurate reconstruction of the forward exact noise vector $\epsilon$, the model tends to overfit the denoising trajectories, resulting in lower reconstruction loss for members compared to non-members. Because DDPMs are too time-consuming for deployment (as they require thousands of denoising steps (Ho et al., 2020)), to reduce the number of denoising steps, more effective consistency-based models have been proposed to learn a globally consistent denoising trajectory. These efficient methods are achieved by incorporating a KL smoothing term into the objective function (Kim et al., 2023; Song et al., 2023; Karras et al., 2022). However, since the KL smoothing term guides the model toward a consistent denoising trajectory across samples rather than memorizing sample-specific trajectories, it reduces the loss discrepancy between member and non-member data and may weaken the effectiveness of intrinsic-based MIAs.
@@ 3.2 Distillation exacerbates I-MIA failures in generative models
Figure 3: Comparison of I-MIAs on teacher model EDM (Karras et al., 2022) and student model DMD (Yin et al., 2024) using: (a) ReDiffuse (Li et al., 2024) reveals membership signals in EDM via distinct reconstruction and re-noising losses between members and non-members on AFHQv2. (b) Applied to DMD, ReDiffuse fails to separate member from non-member instances. (c) Student outputs show stronger distributional alignment with member data when evaluated as instance sets via MMD (Gretton et al., 2012).
Distilled models for online generation service. Although large-scale generative models can produce high-quality text and images, the billion-scale parameters of LLMs and iterative denoise steps of diffusion models lead to high inference latency (Touvron et al., 2023; Song et al., 2020), presenting challenges for online deployment of generation services. Knowledge distillation (Hinton, 2015) offers a compelling solution, enabling the creation of smaller “student” models that learn to mimic the output of larger “teacher” models and achieve gains in inference efficiency without compromising the generation quality (Hsieh et al., 2023; Gu et al., 2023). Examples abound. In LLMs, methods like MiniLLM (Gu et al., 2023) train student models on prompts paired with the teacher’s generated responses. In diffusion models, distillation efforts (Yin et al., 2024; Luo et al., 2024; Song et al., 2023) aim to reduce the multi-step denoising procedure into a single generation step, drastically speeding up inference with a focus on maintaining image quality (Meng et al., 2023; Duan et al., 2023; Luo et al., 2024). This suggests an evolving paradigm of how large-scale generation services are deployed: end-users are increasingly likely to interact with efficient student models, while the computationally intensive teacher models will be dedicated to training these efficient alternatives offline.
@@ 4.2 Distributional statistics enables reliable membership inference with LLMs
Distributional information alleviates the FPR issue. The shift to distributional analysis also addresses a critical flaw plaguing conventional I-MIAs on LLMs (Zhang et al., 2024). As discussed in Sec. 3.1, I-MIAs applied to LLMs often yield an unacceptably high FPR, since computing FPR reliably requires access to “true” non-member data which, model has never truly seen, even indirectly. Given that LLMs are trained on massive web-scale corpora (Liu et al., 2024b), which may contain paraphrases, summaries, or citations of virtually any public data, guaranteeing such pristine non-exposure is practically impossible. Distributional approaches, however, are more resilient to this issue by focusing on aggregate statistical properties of data, such as token frequency distributions or specific sampling behaviors. Even if models see scattered references to non-member datasets, such exposure proves insufficient to reproduce the complete statistical signal of a dataset (Choi et al., 2025). Unless substantial portions of a dataset appear in training, models cannot exhibit the coherent distributional patterns that distributional MIAs detect, enabling reliable membership inference even in contaminated evaluation.
@@ B.1 Preliminaries: MMD and Deep-kernel MMD
This section briefly summarizes the basic knowledge of MMD (Gretton et al., 2012) and its extension, deep-kernel MMD (Liu et al., 2020). MMD finds common application in areas such as domain adaptation and the evaluation of generative models, where assessing distributional alignment is important. In this paper, these concepts are used in quantifying the distributional differences in Section 4 and Section 5. We refer interested readers to the original papers for complete details therein.
Maximum Mean Discrepancy (MMD), proposed by  Gretton et al. (2012), is a statistical tool for measuring the distance between two Borel probability measures, say ${\mathbb{P}}$ and ${\mathbb{Q}}$, defined on a separable metric space ${\mathcal{X}}\subseteq{\mathbb{R}}^{d}$. Consider independent random variables $X,X^{\prime}\sim{\mathbb{P}}$ and $Y,Y^{\prime}\sim{\mathbb{Q}}$. The squared MMD between ${\mathbb{P}}$ and ${\mathbb{Q}}$ in a Reproducing Kernel Hilbert Space ${\mathbb{H}}_{k}$, induced by a kernel function $k:{\mathcal{X}}\times{\mathcal{X}}\to{\mathbb{R}}$, is defined as:
@@
If $k$ is a characteristic kernel (e.g., Gaussian), then $\operatorname{MMD}^{2}(\mathbb{P},\mathbb{Q};k)=0$ if and only if ${\mathbb{P}}={\mathbb{Q}}$.
@@
Deep-kernel MMD. Traditional MMD uses a fixed, pre-defined kernel, which may lead to limited expressiveness when the kernel is not suitable for the task at hand. To address this, Liu et al. (2020) propose to learn a task-relevant representation $\theta_{\omega}:{\mathcal{X}}\to{\mathcal{Z}}$ using a neural network parameterized by $w$. The MMD can then be computed in this learned feature space ${\mathcal{Z}}$. As such, the goal of deep-kernel MMD is to find a representation $\theta_{\omega}$ that maximizes the MMD, thereby increasing the test power to detect differences between ${\mathbb{P}}$ and ${\mathbb{Q}}$.

### v4

@@ 1 Introduction
The success of I-MIAs mainly relies on a phenomenon in training such large-scale generative models: well-trained models tend to overfit to their training set [^1] Note that, for I-MIAs used in classifiers, overfitting is not the only way to help distinguish between members and non-members. For example, Carlini et al. (2022) successfully identified member data by analyzing model behavior signals, such as the output confidence distribution and gradient norms, even when a classifier does not overfit the data. Moreover, it does not mean that the overfitting is a golden rule that always makes I-MIAs successful. When the large-scale models are trained with extremely large datasets (e.g., ChatGPT), I-MIAs could be meaningless (Zhang et al., 2025a). We will discuss this in detail in Section 3.1., exhibiting different behaviors between training instances (i.e., called members) and test instances (i.e., called non-members) (Yeom et al., 2018), which is also known as the memorization effect (Yeom et al., 2018). Based on this phenomenon, many I-MIAs consider using training objectives as criteria to distinguish between members and non-members. For example, for diffusion models, Duan et al. (2023) use estimation errors during the denoising process as the member-detection criterion; for large language models, Ye et al. (2024) and Shejwalkar and Houmansadr (2021) use log-likelihood for the same purpose. Recent studies (Carlini et al., 2023; Hu and Pang, 2023; Duan et al., 2023) have investigated these criteria extensively, hypothesizing that generative models can memorize their training instances.
@@
The failure of I-MIA raises a question: Should we discard the distilled generative models when we care about privacy? The answer is negative, as we discover a distribution-based memory chain between a student-teacher pair: student-generated data exhibit a significantly stronger distributional alignment with teacher’s members than non-members, making it possible to determine if a student model has knowledge from unauthorized data. Specifically, in Sec. 4.1, we use maximum mean discrepancy (MMD) (Gretton et al., 2012) to compare student-generated samples with both teacher’s member and non-member data distributions (Fig. 3(c)). Through repeated random subset samplings, we observe a consistent statistical pattern–distances to non-member data concentrate at higher values than to member data, suggesting that the student preserves statistical signals exhibiting stronger alignment with the teacher’s member distribution than non-member distributions, despite the failure of I-MIAs. Thus, to reliably audit the privacy violation of distilled generative models, we posit the following statement in the field of generative models:
@@ 2 Preliminary: MIAs for Generative Models
We first introduce some basic concepts and relevant literature in this section. For readers who are familiar with MIAs for generative models, it is safe to skip this section. MIAs evaluate whether a specific instance is used during model training (Han et al., 2025). Let $\mathcal{X}$} be a data space and ${\mathcal{D}}^{\rm mem}\subset{\mathcal{X}}$ be a member set used to train a generative model $G:{\mathcal{Z}}\to{\mathcal{X}}$ that transforms noises sampled from a latent distribution ${\mathbf{z}}\sim p({\mathbf{z}})$ into synthetic data ${\mathbf{x}}=f_{g}({\mathbf{z}})\in{\mathcal{X}}$. Given a query sample ${\bm{x}}_{q}\in{\mathcal{X}}$, an MIA constructs a binary classifier ${\mathcal{A}}:{\mathcal{X}}\times G\to\{0,1\}$ that predicts the membership attribution of ${\mathbf{x}}_{q}$ as ${\mathcal{A}}({\bm{x}}_{q},G)=1$ if ${\bm{x}}_{q}\in{\mathcal{D}}^{\rm mem}$, and 0 otherwise. Existing MIAs use instance-level scores to distinguish members from non-members, a strategy we term instance-level MIA (I-MIA), including reference-based and intrinsic-based methods that are introduced below.
@@ 2.1 Brief introduction for instance-level MIAs
where ${\mathcal{D}}^{\rm mem}_{i}$ denotes the training dataset of the $i$-th reference model. The membership inference decision is then based on the difference between the target model $\phi(G,{\bm{x}})$ and these groups. For example, ${\mathcal{A}}({\bm{x}},G)=1$ if a difference metric $\Delta({\bm{x}},G)>\tau$ and 0 otherwise, where $\Delta({\bm{x}},G)\triangleq\mathrm{sim}(\phi(G,{\bm{x}}),{\mathcal{M}}_{1})-\mathrm{sim}(\phi(G,{\bm{x}}),{\mathcal{M}}_{0})$, $\mathrm{sim}(\cdot,\cdot)$ is a function to measure the similarity between two models, and $\tau$ is the decision threshold. However, this approach faces practical infeasibility for computationally intensive generative models like diffusion models, as training $n\geq 1$ reference models would increase the attack cost.
@@ 2.2 How I-MIAs are applied to large-scale generative models?
Large-scale generative model in text domain—LLMs (Guo et al., 2025; Touvron et al., 2023) are auto-regressive transformers (Tang et al., 2024; Chi et al., 2024) trained on massive text corpora to model the log-likelihood of each next token given its preceding context. Several MIA strategies have been developed for LLMs. PCA (Ye et al., 2024) detects membership by comparing the target input with a synthetic version generated by word-swapping. A large log-likelihood gap indicates membership. The Min-K% (Shokri et al., 2017) computes the average log-likelihood of the $k$ tokens with the least confidence in a given input. Typically, member data exhibit higher log-likelihoods for these tokens compared to non-members. Recall (Xie et al., 2024) introduces a non-member prefix to condition the model and quantifies the resulting change in log-likelihood. If the log-likelihood changes significantly, the data is inferred to be a member sample.
Large-scale generative model in vision domain—Diffusion models (Ho et al., 2020; Song et al., 2021), which operate through a forward process $q({\bm{x}}_{t}|{\bm{x}}_{0})$ that progressively adds Gaussian noise $\epsilon\sim{\mathcal{N}}(0,{\mathbf{I}})$ to the input ${\bm{x}}_{0}$ and a reverse noise removal process $p_{G}({\bm{x}}_{0}|{\bm{x}}_{t})$ that reconstructs the original ${\bm{x}}_{0}$ over time steps $t\in[0,T]$. Several attacks have been proposed in this context and have been evaluated on a class of diffusion models, the denoising diffusion probabilistic model (DDPM) (Ho et al., 2020). According to DDPM objective loss $\mathcal{L}_{\text{MSE}}=\left\|\epsilon_{G}(x_{t},t)-\epsilon\right\|^{2}$, SecMI (Duan et al., 2023) performs an attack by monitoring this loss across different time-steps $t$. ReDiffuse (Li et al., 2025) investigates reconstruction stability under noise perturbations, based on the observation that member data tend to yield more consistent reconstructions. GSA (Pang et al., 2023) examines gradient dynamics during model retraining and finds that member data typically induces smaller parameter updates.
@@ 3.1 Large-scale generative models challenge the effectiveness of I-MIAs
I-MIAs face challenges when attacking LLMs. Although I-MIA remains widely used for large-scale generative models, it is unreliable, specifically for LLMs (Dong et al., 2024; Ye et al., 2024). This is because extensive training on massive corpora and substantial model capacity of LLMs would lead to similar behaviors between individual members and non-members exploited by I-MIAs (Ye et al., 2024). When processing an input, LLMs consistently produce high-confidence outputs regardless of whether it is part of training data, reducing the discriminative power of instance-level metrics (Dong et al., 2024). Moreover, Zhang et al. (2025a) argues that I-MIAs on LLMs suffer from unboundable false positive rates (FPR)—a critical metric for evaluating the validity of MIAs when used as evidence to allege the use of unauthorized data (Carlini et al., 2022; Zhang et al., 2025a). The reason is that LLMs are trained with massive web-scale corpora, such that there are no ‘true’ non-members (as LLMs indirectly learn too much data (Zhang et al., 2025a)).
I-MIAs face challenges when attacking diffusion models. In the case of generative models in the visual domain—diffusion models, existing MIAs have successfully identified members in non-distilled settings (Duan et al., 2023; Li et al., 2025). However, there are still issues for both reference-based MIAs and intrinsic-based MIAs. First, reference-based MIAs are impractical for diffusion models due to (1) the high cost of training a reference model (Ho et al., 2020), and (2) the impossibility of recovering the original input noise, which is sampled randomly during training (Ho et al., 2020; Kim et al., 2024).
@@
Second, several intrinsic-based MIAs have shown promising results on DDPMs (Ho et al., 2020). Since the DDPM objective (described in Sec. 2.2) encourages accurate reconstruction of the forward exact noise vector $\epsilon$, the model tends to overfit the denoising trajectories, resulting in lower reconstruction loss for members compared to non-members. Because DDPMs are too time-consuming for deployment (as they require thousands of denoising steps (Ho et al., 2020)), to reduce the number of denoising steps, more effective consistency-based models have been proposed to learn a globally consistent denoising trajectory. These efficient methods are achieved by incorporating a KL smoothing term into the objective function (Kim et al., 2024; Song et al., 2023; Karras et al., 2022). However, since the KL smoothing term guides the model toward a consistent denoising trajectory across samples rather than memorizing sample-specific trajectories, it reduces the loss discrepancy between member and non-member data and may weaken the effectiveness of intrinsic-based MIAs.
@@ 3.2 Distillation exacerbates I-MIA failures in generative models
Figure 3: Comparison of I-MIAs on teacher model EDM (Karras et al., 2022) and student model DMD (Yin et al., 2024) using: (a) ReDiffuse (Li et al., 2025) reveals membership signals in EDM via distinct reconstruction and re-noising losses between members and non-members on AFHQv2. (b) Applied to DMD, ReDiffuse fails to separate member from non-member instances. (c) Student outputs show stronger distributional alignment with member data when evaluated as instance sets via MMD (Gretton et al., 2012).
Distilled models for online generation service. Although large-scale generative models can produce high-quality text and images, the billion-scale parameters of LLMs and iterative denoise steps of diffusion models lead to high inference latency (Touvron et al., 2023; Song et al., 2021), presenting challenges for online deployment of generation services. Knowledge distillation (Hinton, 2015) offers a compelling solution, enabling the creation of smaller “student” models that learn to mimic the output of larger “teacher” models and achieve gains in inference efficiency without compromising the generation quality (Hsieh et al., 2023; Gu et al., 2023). Examples abound. In LLMs, methods like MiniLLM (Gu et al., 2023) train student models on prompts paired with the teacher’s generated responses. In diffusion models, distillation efforts (Yin et al., 2024; Luo et al., 2024; Song et al., 2023) aim to reduce the multi-step denoising procedure into a single generation step, drastically speeding up inference with a focus on maintaining image quality (Meng et al., 2023; Duan et al., 2023; Luo et al., 2024). This suggests an evolving paradigm of how large-scale generation services are deployed: end-users are increasingly likely to interact with efficient student models, while the computationally intensive teacher models will be dedicated to training these efficient alternatives offline.
@@ 4.2 Distributional statistics enables reliable membership inference with LLMs
Distributional information alleviates the FPR issue. The shift to distributional analysis also addresses a critical flaw plaguing conventional I-MIAs on LLMs (Zhang et al., 2025a). As discussed in Sec. 3.1, I-MIAs applied to LLMs often yield an unacceptably high FPR, since computing FPR reliably requires access to “true” non-member data which, model has never truly seen, even indirectly. Given that LLMs are trained on massive web-scale corpora (Liu et al., 2024b), which may contain paraphrases, summaries, or citations of virtually any public data, guaranteeing such pristine non-exposure is practically impossible. Distributional approaches, however, are more resilient to this issue by focusing on aggregate statistical properties of data, such as token frequency distributions or specific sampling behaviors. Even if models see scattered references to non-member datasets, such exposure proves insufficient to reproduce the complete statistical signal of a dataset (Choi et al., 2025). Unless substantial portions of a dataset appear in training, models cannot exhibit the coherent distributional patterns that distributional MIAs detect, enabling reliable membership inference even in contaminated evaluation.
@@ B.1 Preliminaries: MMD and Deep-kernel MMD
This section briefly summarizes the basic knowledge of MMD (Gretton et al., 2012) and its extension, deep-kernel MMD (Liu et al., 2020). MMD finds common application in areas such as domain adaptation (Jiang et al., 2023; Chi et al., 2021; Zheng et al., 2023) and the evaluation of generative models, where assessing whether two or more domains are distributionally aligned (Zhong et al., 2024; Liu et al., 2021; Ye and Yao, 2022; Ye et al., 2023), is essential. MMD and related techniques have been extensively adopted in real-world applications, including healthcare (Zhong et al., 2024; Guo et al., 2022; Jiang et al., 2016) and defense against adversarial attacks (Zhang et al., 2025b, 2023). In this paper, these concepts are used in quantifying the distributional differences in Section 4 and Section 5. We refer interested readers to the original papers for complete details therein.
Maximum Mean Discrepancy (MMD), proposed by Gretton et al. (2012), is a statistical tool for measuring the distance between two Borel probability measures, say $\mathbb{P}$} and $\mathbb{Q}$}, defined on a separable metric space ${\mathcal{X}}\subseteq{\mathbb{R}}^{d}$. Consider independent random variables $X,X^{\prime}\sim{\mathbb{P}}$ and $Y,Y^{\prime}\sim{\mathbb{Q}}$. The squared MMD between $\mathbb{P}$} and $\mathbb{Q}$} in a Reproducing Kernel Hilbert Space ${\mathcal{H}}_{k}$, induced by a kernel function $k:{\mathcal{X}}\times{\mathcal{X}}\to{\mathbb{R}}$, is defined as:
@@
If $k$ is a characteristic kernel (e.g., Gaussian), then $\operatorname{MMD}^{2}(\mathbb{P},\mathbb{Q};k)=0$ if and only if $\mathbb{P}$$\mathbb{Q}$}.
@@
Deep-kernel MMD. Traditional MMD uses a fixed, pre-defined kernel, which may lead to limited expressiveness when the kernel is not suitable for the task at hand. To address this, Liu et al. (2020, 2021) propose to learn a task-relevant representation $\theta_{\omega}:{\mathcal{X}}\to{\mathcal{Z}}$ using a neural network parameterized by $w$. The MMD can then be computed in this learned feature space $\mathcal{Z}$}. As such, the goal of deep-kernel MMD is to find a representation $\theta_{\omega}$ that maximizes the MMD, thereby increasing the test power to detect differences between $\mathbb{P}$} and $\mathbb{Q}$}.